# Cross-lingual Human-Preference Alignment for Neural Machine Translation with Direct Quality Optimization

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) and derivative techniques like Direct Preference Optimization (DPO) are task-alignment algorithms used to repurpose general, foundational models for specific tasks. We show that applying task-alignment to neural machine translation (NMT) addresses an existing task–data mismatch in NMT, leading to improvements across all languages of a multilingual model, even when task-alignment is only applied to a subset of those languages. We do so by introducing Direct Quality Optimization (DQO), a variant of DPO leveraging a pre-trained translation quality estimation model as a proxy for human preferences, and verify the improvements with both automatic metrics and human evaluation.
@@ 1 Introduction
    Improvements in held-out languages are not limited to general behaviors required by the translation task (e.g., avoiding source language fragments, translation additions and omissions), but include language-specific linguistic features not seen in the DQO alignment data, such as transliteration of named entities in Latvian.
While we attribute much of the performance in held-out languages to transfer learning of general behaviors required by the translation task (such as avoiding source language fragments and translation additions, omissions, or inconsistencies), the language-specific improvements in held-out languages cannot be explained by transfer learning.
@@ 2.2 Training Data Mismatch
2.2.1 Web Data Mining
A large portion of parallel data is mined from massive collections of web documents, using automated methods to detect and align source and target language segments – the popular datasets ParaCrawl (Bañón et al., 2020) and CCMatrix (Schwenk et al., 2021b), for example. This process may capture human translations, text written independently in both the source and target languages on the same topic, or the output of other MT models.
One prominent cause of a task–data mismatch in automatically aligned sentence pairs is semantic misalignment. Kreutzer et al. (2022) found semantic misalignment in 15% (ParaCrawl) and 32% (CCMatrix) of sentence pairs as part of a manual quality audit.
@@ 2.2.1 Web Data Mining
Unfortunately, slight semantic misalignments of source and target are both more prevalent and much more difficult for state-of-the-art filtering systems to detect Meng et al. (2024). These may include subtle yet significant differences in meaning, factual differences in numbers or names, additions and omissions, and the accompanying losses in translation adequacy.
In addition, these segments often still contain useful information that may help the model learn Meng et al. (2024).
2.2.2 Accidental Inclusion of Machine Translated Content
Web data may also include the outputs of other machine translation models, including neural, statistical and dictionary-based methods of varying quality. The impact of training on low quality machine translations is clear, however even good NMT systems’ outputs differ significantly enough from natural text that classifiers can be trained to detect machine translated text with high accuracy – and even predict which machine translation system was used to translate a given text (La Morgia et al., 2023).
@@ 2.2.2 Accidental Inclusion of Machine Translated Content
2.2.3 Translator Skill Level
Another source of task–data mismatch in human translations is the fact that human translators differ in skill level Albir (2017). This implies that not all human translations will be equally preferred by humans.
@@ 2.2.3 Translator Skill Level
2.2.4 Translationese
Another significant issue is a phenomenon known as translationese, the observation that human-translated texts in a given language differ in distribution from texts written independently in that language. Specifically, translated text shows signs of interference from the source language’s grammar, word order and word choice, as well as source language independent effects of the process of translation itself, such as simplification and avoidance of unique language features (Koppel and Ordan, 2011; Laviosa, 1998; Tirkkonen-Condit, 2004).
These effects are significant enough that classification models can distinguish translated and original text with high accuracy (Baroni and Bernardini, 2005; Sominsky and Wintner, 2019), as well as identifying the source language of the text (Koppel and Ordan, 2011).
As humans show a consistent preference for translations closer to the distribution of original text rather than translationese (Riley et al., 2020; Freitag et al., 2022b), this creates an inherent task–data mismatch for training data translated in the source–target direction.
2.2.5 Source–Target Domain Mismatch
Translation pairs in the other direction, target–source, are better aligned with human preference, as the target labels are drawn from the original text distribution rather than from translationese.
Unfortunately, they suffer from another subtle source of task–data mismatch found in human translations: source–target domain mismatch (Shen et al., 2021).
Source–target domain mismatch is the observation that speakers of different languages tend to discuss different topics. For instance, a Cherokee newspaper is likely to report on different topics than an Icelandic newspaper would, and translations of these articles would remain representative of the Cherokee or Icelandic language domains, respectively.
This effect is especially pronounced for low-resource language pairs (Shen et al., 2021).
@@ 3 Human Preference Learning for LLMs
Supervised data showing chat-based dialog between humans and AI assistants was, prior to the wide availability of such agents in the form of LLMs, understandably rare. The only possible method of creating such data was to hire humans to role play as AI assistants – an expensive endeavor that few research teams had the funding or time to undertake. Even with the advent of high quality proprietary and open-source models, which one could sample to create synthetic data, there is a fundamental task–data mismatch: the task is not to imitate an existing AI assistant, but (ideally) to train a new state-of-the-art model.
@@
This is then followed by instruction fine-tuning, a comparatively brief round of supervised learning on human or AI labeled examples of dialogues, which brings the model’s output distribution into the general neighborhood of desired behavior. Finally human preference learning, using actual human rankings aligns the model with the desired task: producing human-preferred responses to questions and dialog, while remaining helpful and harmless (Bai et al., 2022).
@@ 4 Direct Quality Optimization for NMT
Because of its stability and ease of use, we select DPO as the basis for our experiments with human preference learning as a form of task alignment. As a proxy for human preferences, we use the CometKiwi quality estimation model to score and compare multiple translations of a given source Rei et al. (2022b). CometKiwi is highly multilingual and has been shown to correlate well with human preference Kocmi et al. (2024).
@@
Table 1: Target languages supported by the NVIDIA Megatron model. The category “Other” contains all languages that are the only supported representative of their language family. The languages on which we apply task alignment are in boldface.
@@
Any improvements in those languages must either apply to all languages (such as avoiding omissions or additions), or are language specific, and can only have come from previously unused latent knowledge from supervised training.
@@
As the seed dataset from which to draw source sentences for human preference learning, we use the source side of a mixture of publicly available English–German MT datasets (listed in Appendix A.3), with the goal of covering a wide range of source domains.
@@
This results in slightly under 8000 preference pairs (occasionally the maximum difference in COMET score between a segment’s highest and lowest scoring sampled translations is less than $\varepsilon$, in which case we do not produce a preference pair), we run DPO training with a batch size of 8192 tokens (counting source, chosen and rejected tokens), a learning rate of $1\mathrm{e}{-6}$ and $\beta=0.5$. A complete list of hyperparameters can be found in Appendix 7.
At this point, we train on the preference pairs using standard DPO for 8 epochs, after which we sample a fresh set of source segments from the seed dataset, sample translations from the policy model, create a new set of preference pairs, and begin the training again. This helps ensure that the preference pairs used are relevant to the policy model throughout training.
In total, we perform 6 such rounds of DPO training. We call this end-to-end process Direct Quality Optimization (DQO), detailed formally in Figure 1.
Parameters: preference relation $\succ$, number of rounds $n$, epochs per round $m$, epoch size $d$, learning rate $\alpha$, DPO regularization $\beta$, number of sampled translations per source $k$
@@
[…]
This can be viewed as a batched online version of DPO, as the updates are performed on batches of data sampled from the policy model. Initial experiments showed that performance gains rapidly plateaued under standard DPO with a static dataset of preference pairs.
Model Lang. FLORES+ NTREX
BLEURT COMET CometKiwi BLEU BLEURT COMET CometKiwi BLEU
Baseline All 0.7732 0.8787 0.8451 34.33 0.7127 0.8414 0.8169 30.76
DQO All 0.7928 0.8923 0.8585 35.45 0.7354 0.8593 0.8344 31.68
Baseline $\mathcal{T}$ 0.7329 0.8467 0.8334 34.88 0.6795 0.8120 0.8070 33.19
DQO $\mathcal{T}$ 0.7498 0.8615 0.8478 35.51 0.7009 0.8309 0.8256 33.74
Baseline $\mathcal{T}^{c}$ 0.7812 0.8851 0.8475 34.22 0.7193 0.8473 0.8189 30.27
DQO $\mathcal{T}^{c}$ 0.8014 0.8985 0.8606 35.43 0.7422 0.8650 0.8362 31.27
Baseline $\mathcal{R}\cap\mathcal{T}^{c}$ 0.7909 0.8864 0.8510 36.58 0.7297 0.8478 0.8213 33.44
DQO $\mathcal{R}\cap\mathcal{T}^{c}$ 0.8080 0.8979 0.8624 37.61 0.7508 0.8648 0.8377 34.58
Baseline $\mathcal{R}^{c}$ 0.7689 0.8833 0.8431 31.22 0.7061 0.8465 0.8158 26.23
DQO $\mathcal{R}^{c}$ 0.7930 0.8992 0.8583 32.66 0.7313 0.8652 0.8342 27.05
Table 3: Evaluation metrics on FLORES+ and NTREX with the NVIDIA Megatron EN-X model, before and after task-alignment using DQO. Results are shown for relevant groupings of the 30 target languages: all languages, languages used in DQO ($\mathcal{T}$), languages not used in DQO ($\mathcal{T}^{c}$), languages not used in DQO, but related to those used ($\mathcal{R}\cap\mathcal{T}^{c}$), and languages neither used nor related to the languages used ($\mathcal{R}^{c}$).
@@ 5.1 Automatic Quality Metrics
We use corpus-level sacreBLEU [^3] Signature: nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.4.0. For JA and ZH, we additionally use the mecab-ja and mecab-zh tokenizers. Post (2018) as well as three neural evaluation models:
  1. 1.
    Reference-free CometKiwi (Rei et al., 2022b)
  2. 2.
    Reference-based COMET (Rei et al., 2022a)
  3. 3.
    BLEURT (Sellam et al., 2020)
@@
Figure 2: Changes in BLEURT on FLORES+ and NTREX with the NVIDIA Megatron EN-X model, before and after task alignment with DQO. Languages used in DQO are bolded.
Results are reported in Table 3 and Figure 2. We find that DPO task alignment increases all three neural quality metrics on both datasets for each of the 30 target languages supported by the Megatron EN-X model.
BLEU scores increased for all languages on both datasets, with the exception of Hindi, which decreased by $0.70$ BLEU on NTREX and $1.12$ BLEU on FLORES+, despite showing improvements on the three neural metrics, like all other languages. The exact cause of this exception is unclear, especially as Hindi was one of the five languages used for DPO task alignment.
@@ 5.2 Training Data Perplexity
In order to confirm the existence of a task–data mismatch, we examine how DQO affects the model’s perplexity on the training data. As we do not have access to the training data used for the NVIDIA Megatron English-Many model, we repeat the above experiment with a proprietary encoder–decoder model trained on publicly available English-to-German data using the NVIDIA NeMo framework (Kuchaiev et al., 2019). For the full list of training datasets, see Appendix A.3.
The model architecture is similar to the Megatron model, with both following the deep encoder, shallow decoder recipe suggested by Kasai et al. (2021), but is significantly larger, with an embedding size of 2048, a feed-forward width of 8192, 21 encoder layers and 2 decoder layers, and a 32768 token vocabulary, resulting in a total of 1.3B parameters.
We apply DQO to this model as with the Megatron model, however using only English–German preference pairs.
After applying DQO, we see large improvements in CometKiwi and COMET for a variety of evaluation datasets, confirming that DQO worked as expected.
Source The first economic reforms were made under the leadership of Deng Xiaoping.
Baseline Pirmās ekonomiskās reformas tika veiktas Deng Xiaoping vadībā.
DQO Pirmās ekonomiskās reformas tika veiktas Dena Sjaopina vadībā.
@@
Source However, a source with knowledge of the investigation told the Minneapolis Star-Tribune that it was Louis Jourdain, 16-year old son of Red Lake Tribal Chairman Floyd Jourdain.
Baseline Tomēr avots, kas zina par izmeklēšanu, Minneapolis Star-Tribune pastāstīja, ka tas bija Louis Jourdain, 16 gadus vecs Red Lake cilts priekšsēdētāja Floida Jourdaina dēls.
DQO Tomēr avots, kas ir informēts par izmeklēšanu, Minneapolis Star-Tribune pastāstīja, ka tas bija Luiss Džordēns, 16 gadus vecs Red Lake cilts priekšsēdētāja Floida Džordēna dēls.
Source King Sejong was the fourth king of the Joseon Dynasty and is one of the most highly regarded.
Baseline King Sejong bija ceturtais karalis no Joseon dinastijas un ir viens no visvairāk augstu vērtē.
DQO Karalis Sedžons bija ceturtais Džosona dinastijas karalis un ir viens no visaugstāk novērtētajiem.
Table 4: Examples of translations into Latvian from the FLORES+ data set before and after DQO. Names are bolded to highlight the DQO model’s increased ability to consistently transliterate names into Latvian orthography. Names that are incorrectly transliterated are in italics.
The arithmetic mean of perplexity over a random sample of 1 million segments from the training data increased from $7.219$ (baseline model) to $9.435$ (DQO), confirming that the improvements in preference are not due to increased ability to model the training data.

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) and derivative techniques like Direct Preference Optimization (DPO) are task-alignment algorithms used to repurpose general, foundational models for specific tasks. We show that applying task-alignment to neural machine translation (NMT) addresses an existing task–data mismatch in NMT, leading to improvements across all languages of a multilingual model, even when task-alignment is only applied to a subset of those languages. We do so by introducing Direct Quality Optimization (DQO), a variant of DPO leveraging a pre-trained translation quality estimation model as a proxy for human preferences, and verify the improvements with both automatic metrics and through human evaluation.
Cross-lingual Human-Preference Alignment for
Neural Machine Translation with Direct Quality Optimization
Kaden Uhlig  and Joern Wuebker  and Raphael Reinauer  and John DeNero LILT {kaden.uhlig,joern,raphael.reinauer,john}@lilt.com
@@ 1 Introduction
    Improvements in held-out languages are not limited to general behaviors required by the translation task (e.g. avoiding omissions), but include improving language-specific linguistic features not seen in the DQO alignment data, such as correctly transliterating named entities in Latvian.
While we attribute much of the performance in held-out languages to transfer learning of general behaviors required by the translation task (e.g. avoiding omissions), the language-specific improvements in held-out languages cannot be explained by transfer learning.
@@ 2.2 Training Data Mismatch
Web Data Mining. A large portion of parallel data is mined from massive collections of web documents, using automated methods to align source and target language segments – e.g. the ParaCrawl (Bañón et al., 2020) and CCMatrix (Schwenk et al., 2021b) datasets. This process may capture human translations, text written independently in both the source and target languages on the same topic, or the output of other MT models. One prominent cause of task–data mismatch in automatically aligned sentence pairs is semantic misalignment. Kreutzer et al. (2022) found semantic misalignment in 15% (ParaCrawl) and 32% (CCMatrix) of sentence pairs in a manual quality audit.
@@
Unfortunately, slight semantic misalignments of source and target are both more prevalent and much more difficult for state-of-the-art filtering systems to detect Meng et al. (2024). These may include subtle yet significant differences in meaning, factual differences in numbers or names, additions and omissions, and the accompanying losses in translation adequacy. In addition, these segments often still contain useful information that may help the model learn Meng et al. (2024).
Accidental Inclusion of Machine Translated Content. Web data may also include the outputs of other machine translation models, including neural, statistical and dictionary-based methods of varying quality. The impact of training on low quality machine translations is clear, however even good NMT systems’ outputs differ significantly enough from natural text that classifiers can detect machine translated text with high accuracy – and even predict which machine translation system was used to translate a given text (La Morgia et al., 2023).
@@
Translator Skill Level. Another source of task–data mismatch in human translations is the fact that translators differ in skill level Albir (2017). This implies that not all human translations will be equally preferred by humans.
@@
Model Lang. FLORES+ devtest NTREX
BLEURT COMET22 CometKiwi22 BLEU BLEURT COMET22 CometKiwi22 BLEU
Baseline All 0.7614 0.8741 0.8387 34.19 0.7016 0.8359 0.8099 30.31
DQO All 0.7790 0.8873 0.8508 35.31 0.7212 0.8525 0.8255 31.21
Baseline $\mathcal{T}$ 0.7231 0.8417 0.8272 34.50 0.6677 0.8040 0.7979 32.62
DQO $\mathcal{T}$ 0.7381 0.8559 0.8401 35.34 0.6854 0.8209 0.8137 33.16
Baseline $\mathcal{T}^{c}$ 0.7691 0.8805 0.8410 34.13 0.7084 0.8423 0.8123 29.85
DQO $\mathcal{T}^{c}$ 0.7872 0.8935 0.8529 35.30 0.7284 0.8588 0.8278 30.82
Baseline $\mathcal{R}\cap\mathcal{T}^{c}$ 0.7802 0.8820 0.8447 36.46 0.7202 0.8432 0.8154 33.01
DQO $\mathcal{R}\cap\mathcal{T}^{c}$ 0.7967 0.8936 0.8557 37.54 0.7391 0.8593 0.8307 34.13
Baseline $\mathcal{R}^{c}$ 0.7549 0.8787 0.8364 31.17 0.6934 0.8413 0.8084 25.84
DQO $\mathcal{R}^{c}$ 0.7751 0.8934 0.8493 32.46 0.7147 0.8581 0.8242 26.61
Table 1: Evaluation metrics on FLORES+ devtest and NTREX with the NVIDIA Megatron EN-X model, before and after task-alignment using DQO. Results are shown for relevant groupings of the 30 target languages: all languages, languages used in DQO ($\mathcal{T}$), languages not used in DQO ($\mathcal{T}^{c}$), languages not used in DQO, but related to those used in DQO ($\mathcal{R}\cap\mathcal{T}^{c}$), and languages neither used nor related to the languages used in DQO ($\mathcal{R}^{c}$).
Translationese. Another common issue is a phenomenon known as translationese, the observation that human-translated text in a given language differ in distribution from text written independently in that language. Specifically, translated text shows signs of interference from the source language’s grammar, word order and word choice, as well as source-language-independent effects of the translation process itself, such as simplification and avoidance of unique language features (Koppel and Ordan, 2011; Laviosa, 1998; Tirkkonen-Condit, 2004). These effects are significant enough that classification models can distinguish translated and original text with high accuracy (Baroni and Bernardini, 2005; Sominsky and Wintner, 2019), as well as identifying the source language of the text (Koppel and Ordan, 2011). As humans show a consistent preference for translations closer to the distribution of original text rather than translationese (Riley et al., 2020; Freitag et al., 2022b), this creates an inherent task–data mismatch for training data translated in the source–target direction.
Source–Target Domain Mismatch. Translation pairs in the other direction, target–source, are better aligned with human preference, as the target labels are drawn from the original text distribution rather than from translationese. Unfortunately, they suffer from another subtle source of task–data mismatch found in human translations: Source–target domain mismatch (Shen et al., 2021) is the observation that speakers of different languages tend to discuss different topics. For instance, a Cherokee newspaper is likely to report on different topics than an Icelandic newspaper would, and translations of these topics would remain representative of the Cherokee domain. This effect is especially pronounced for low-resource language pairs (Shen et al., 2021).
@@ 3 Human Preference Learning for LLMs
Supervised data showing chat-based dialog between humans and AI assistants was, prior to the wide availability of such agents in the form of LLMs, understandably rare. Even with the advent of high quality proprietary and open-source models, which one could sample to create synthetic data, there is a fundamental task–data mismatch: the task is not to imitate an existing AI assistant, but (ideally) to train a new state-of-the-art model.
@@
This is then followed by instruction fine-tuning, a comparatively brief round of supervised learning on human- or AI-labeled examples of dialogues, which brings the model’s output distribution into the general neighborhood of desired behavior. Finally human preference learning, using actual human rankings aligns the model with the desired task: producing human-preferred responses to questions and dialog, while remaining helpful and harmless (Bai et al., 2022).
@@ 4 Direct Quality Optimization for NMT
Because of its stability and ease of use, we select DPO as the basis for our experiments with human preference learning as a form of task alignment for NMT. As a proxy for human preferences, we use the CometKiwi quality estimation model to score and compare multiple translations of a given source Rei et al. (2022b). CometKiwi is highly multilingual and has been shown to correlate well with human preference Kocmi et al. (2024).
@@
Table 2: Target languages supported by the NVIDIA Megatron En-X model. The category “Other” contains all languages that are the only supported representative of their language family. The languages on which we apply task alignment are denoted in boldface.
@@
Any improvements in those languages must either apply to all languages (such as avoiding omissions or additions), or are language specific, and can only have come from previously unused latent knowledge acquired during supervised training.
@@
As the seed dataset from which to draw source sentences for human preference learning, we use the source side of a mixture of publicly available English–German MT datasets (see Appendix A.3).
@@
This results in slightly under 8000 preference pairs (occasionally the maximum difference in COMET score between a segment’s highest and lowest scoring sampled translations is less than $\varepsilon$, in which case we do not produce a preference pair), we run DPO training with a batch size of 8192 tokens (counting source, chosen and rejected tokens), a learning rate of $1\mathrm{e}{-6}$ and $\beta=0.5$. See Appendix 7 for a full list of hyperparameters.
At this point, we train on the preference pairs using standard DPO for 8 epochs, after which we sample a fresh set of source segments from the seed dataset, sample translations from the policy model, create a new set of preference pairs, and begin the training again. This resampling process helps ensure that the preference pairs remain relevant to the policy model during training, and leads to substantial performance improvements. In total, we perform 5 rounds of DPO training. We call this end-to-end process Direct Quality Optimization (DQO), detailed formally in Algorithm 1.
Parameters: preference relation $\succ$, number of rounds $n$, epochs per round $m$, epoch size $d$, learning rate $\alpha$, DPO regularization $\beta$, sampled translations per source $k$
@@
Figure 1: Direct Quality Optimization (DQO). Greedy ($\!{}_{\pi}(x)$ is the translation of $x$ produced with greedy search and the model $\pi$. DPO refers to Direct Preference Optimization – for full implementation details see Rafailov et al. (2023).
DQO can be viewed as a batched online version of DPO, as the updates are performed on batches of data sampled from the policy model.
@@ 5.1 Automatic Quality Metrics
We use corpus-level sacreBLEU [^3] Signature: nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.4.0. For JA and ZH, we additionally use the mecab-ja and mecab-zh tokenizers. Post (2018) as well as three neural evaluation models: Reference-free CometKiwi (Rei et al., 2022b), reference-based COMET (Rei et al., 2022a), and BLEURT (Sellam et al., 2020).
@@
Figure 2: Changes in BLEURT on FLORES+ devtest with the NVIDIA Megatron EN-X model, before and after task alignment with DQO. Languages used in DQO are bolded.
Results are reported in Table 1 and Figure 2. We find that DPO task alignment increases all three neural quality metrics on both datasets for each of the 30 target languages. BLEU scores increased for all languages on both datasets, with the exception of Hindi, which decreased by $0.40$ BLEU on NTREX and $0.4$ BLEU on FLORES+ devtest, despite showing improvements on the three neural metrics, like all other languages.
@@
Figure 3: Mean BLEURT-20 on FLORES+ dev at each round of DQO with the NVIDIA Megatron EN-X model, using either Direct Quality Optimization (DQO) or Supervised Fine-Tuning (RAFT) to update the model.
To ablate the use of DPO as the update step within DQO, we perform a comparative experiment identical to DQO as described in Section 4 and Algorithm 1, but using standard supervised fine-tuning (SFT) on the preferred translation instead of DPO. Note that this becomes essentially equivalent to Reward rAnked Fine-Tuning (RAFT) Dong et al. (2023). Figure 3 shows mean performance for all language pairs through the 5 rounds of DQO. See Appendix A.5 for individual languages.
@@ 5.2 Training Data Perplexity
In order to confirm the existence of a task–data mismatch, we examine DQO’s effect on model perplexity over the training data. As we do not have access to the training data used for the NVIDIA Megatron English-Many model, we repeat the above experiment with a proprietary encoder–decoder model trained on publicly available English-to-German data using the NVIDIA NeMo framework (Kuchaiev et al., 2019) (See Appendix A.3). The model architecture is similar to the Megatron model, and follows the deep encoder, shallow decoder recipe from Kasai et al. (2021), but is larger, with a model width of 2048, a feed-forward width of 8192, 21 encoder layers, 2 decoder layers, and a 32k token vocabulary, totaling 1.3B parameters.
We apply DQO to this model as with the Megatron model, however using only English–German preference pairs. After applying DQO, we see large improvements in CometKiwi and COMET for a variety of evaluation datasets, confirming that DQO worked as expected. The arithmetic mean of perplexity over a random sample of 1 million segments from the training data increased from $7.219$ (baseline model) to $9.435$ (DQO), confirming that the improvements in test data preference correspond to a reduction in the model’s fit to the training corpus.
Source … under the leadership of Deng Xiaoping.
Baseline … tika veiktas Deng Xiaoping vadībā.
DQO … tika veiktas Dena Sjaopina vadībā.
@@
Source … that it was Louis Jourdain, 16-year old son of … Floyd Jourdain.
Baseline … ka tas bija Louis Jourdain, 16 gadus vecs … Floida Jourdaina dēls.
DQO … ka tas bija Luiss Džordēns, 16 gadus vecs … Floida Džordēna dēls.
Source King Sejong was the fourth king of the Joseon Dynasty …
Baseline King Sejong bija ceturtais karalis no Joseon dinastijas …
DQO Karalis Sedžons bija ceturtais Džosona dinastijas karalis …
Table 4: Examples of translations into Latvian from the FLORES+ data set before and after DQO. Names are bolded to highlight the DQO model’s increased ability to consistently transliterate names into Latvian orthography. Names that are incorrectly transliterated are in italics. Sentences are truncated to avoid dataset leakage.